OpenStudy (anonymous):

-1/2 + -7/8 = -4/5 -1 3/8 -1 3/11 -3/5

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2 }+ -\frac{ 7 }{ 8 }\] Change these so your fractions are the same. (Common denominator.) so \[-\frac{ 4 }{ 8 } + -\frac{ 7 }{ 8 } = -\frac{ 11 }{ 8 } or -1\frac{ 3 }{ 8 }\]